Carry is 9 years older than Dan. In 8 years thee sum of their age will be 97. Find the age of each man now.
Answer by addingup(3677)

C = D+9
C+8+D+8 = 97 Substitute for C
D+9+8+D+8 = 97
2D+25 = 97
2D = 72
D = 36
and C = D+9 = 36+9 = 45
----------------------
Check - in 8 years:
45+8+36+8 = 97
53+44 = 97 Add the left and you'll see that the answer is correct.
:
john